OpenStudy (anonymous):

What is the simplified form of 15x^5/9y^7• 45y^3/5x^2 ?

OpenStudy (anonymous):

3x^3 5y^-4

OpenStudy (anonymous):

\[\frac{ 15x^3 }{ y^4 }\]

OpenStudy (anonymous):

\[\frac{ 15x^4 }{ y^3 }\]

OpenStudy (anonymous):

\[\frac{ y^4 }{ 15x^3 }\]

OpenStudy (anonymous):

\[\frac{ y^3 }{ 15x^4 }\]

OpenStudy (anonymous):

those are the options

OpenStudy (anonymous):

It would be 15x^3/y^4